Write the Fourier Series of the function f (t) = | cos (t) | for t defined on the interval
[−π, π].
Answer #1

Solna. Given fct) = [cost] in 1-1,1) f(t) = cost is an Even fn Since Hence, bn=0 We Can Write fat = Cost ) 0<< - Cost, I TACTDATE Now, 다. an flott) Cosnet dat IT 0 Cost Cosnt dot t 7) IT f(- Cost] Cosnate dat T2 T/2 Å f{cos (n+1)* + Cos (n-1)+ dt 2 {Cos ntl Cos II nt 9 o Sin (nT + 1) = Cos & Sin I ) Cos ni g 2 dan Cos nI - 4 (28) nh When ne T Now, when (n=1) Then a = 2 2G DATE 1) It [ I - TT+T 2 2 | a=0 frath =2 to ㅠ y ų E I Corni Cosnt Cosnats T n=2 (02-1) 9 1 688+) = -4 Fcosat + corput ---]